# EvaLoop: Assessing LLM Robustness in Programming from a Self-consistency Perspective

### v1

@@ 1 Introduction
Evaluating the robustness of LLMs in programming often involves adversarial attacks anand2021adversarial ; jha2023codeattack ; nguyen2023adversarial ; chen2023evaluating ; zhang2024attacks ; improta2025enhancing ; wei2023towards ; zaremba2025trading . In these methods, natural language prompts are perturbed at various levels (e.g., character, word, sentence), and the functional correctness of code generated from original versus perturbed prompts is compared to measure performance degradation, regarding as the robustness measurement. Although these adversarial attacks can quantify the robustness, their reliability and generalization face challenges. Our observations indicate that different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed. Specifically, when we applied three adversarial attacks—character, word, and sentence levels respectively—used in recode wei2023towards to the MBPP Plus benchmark liu2024your , and instructed GPT-3.5-Turbo to generate results for both original and adversarial prompts. As a result, GPT-3.5-Turbo only performs less robustly under word-level adversarial attack with synonym replacement. These findings suggest that a single adversarial attack may not provide a holistic view of an LLM’s robustness in programming, and a suite of diverse attacks might be necessary for a comprehensive assessment of a single model. However, this approach complicates the comparative evaluation of robustness across different models. For instance, if one model demonstrates strong robustness against attack X but performs poorly against attack Y, while another model shows the opposite pattern, it becomes challenging to definitively determine which model is inherently more robust overall. This motivates a question: Can we overcome the configuration-dependency of adversarial attacks to achieve a unified evaluation of LLMs’ robustness in programming?
@@ 3 Evaluation
Benchmark Selection. EvaLoop is designed for versatility and seamless integration with existing programming benchmarks. In this study, we utilize two comprehensive benchmark suites, MBPP Plus liu2024your and MXEVAL athiwaratkun2022multi , which offer significant advancements over their predecessors. MBPP Plus enhances the original MBPP austin2021program by applying rigorous quality filtering to programming tasks and expanding the test case corpus 35-fold. MXEVAL extends MBPP’s Python-centric focus to encompass 12 distinct programming languages. For Framework 1 (Generation-Summarization Duality Loop), we evaluate the LLMs’ robustness in programming on 378 distinct tasks from MBPP Plus. The operational loop for this framework is limited to a maximum of 10 loops. For Framework 2 (Cyclical Code Translation Loop), we establish a circular code translation chain: Python → JavaScript → PHP → Ruby → Perl → Python. The four intermediate languages were randomly selected from the non-Python languages available in MXEVAL [^3] Our experimental results showed no significant differences between two such distinct random selections.. This cyclical translation process utilizes a carefully curated subset of 375 tasks from the multilingual MXEVAL.
@@
Examining ASL metric reveals another insight into LLMs’ robustness in programming. GPT-4.1 leads with an ASL of 3.73, closely followed by GPT-4.1-mini (3.72) and o3-mini (3.71), signifying not only strong initial accuracy but also remarkable performance consistency across sequential translation tasks, underscoring their robust nature and the superior. In stark contrast, open-source LLMs generally represent a relatively low ASL value, and LLaMA-3.1-8B records the lowest ASL with 2.37, whose trajectory shows a low baseline followed by continuous, pronounced degradation, highlighting difficulties in maintaining functional integrity. ASL is invaluable for a holistic assessment, moving beyond single-point evaluations to capture sustained operational efficacy and distinguish models with enduring performance integrity from those that falter under the cumulative stress of multi-step, cross-lingual code translations.
@@ 4 Related Work
LLMs’ Robustness in Programming. Code generation has emerged as the predominant practical application of AI in programming environments liang2024large , catalyzing extensive research into the robustness of LLMs in this domain. Empirical assessments of commercial systems reveal concerning vulnerabilities: GitHub Copilot produces inconsistent implementations for semantically equivalent prompts in 46% of cases, with correctness variations reaching 28% between prompt variants mastropaolo2023robustness . Similarly, Codex exhibits substantial sensitivity to minor linguistic perturbations in task specifications shirafuji2023exploring ; zhuo2023robustness . The research community has responded with LLMs’ robustness through diverse adversarial techniques chen2023evaluating ; zhang2024attacks ; jha2023codeattack alongside corresponding defensive strategies improta2025enhancing ; yang2024robustness ; ge2024demonstration . However, different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed, struggling in a unifed evaluation. EvaLoop addresses this critical gap by leveraging the duality between generation and understanding, enabling unified robustness assessment across LLMs with the average number of sustainable loops, i.e., ASL . This novel approach provides a natural measure of LLMs’ robustness in programming.

### v2

@@ 1 Introduction
Evaluating the robustness of LLMs in programming often involves adversarial attacks anand2021adversarial ; jha2023codeattack ; nguyen2023adversarial ; chen2023evaluating ; zhang2024attacks ; improta2025enhancing ; wei2023towards ; zaremba2025trading ; bavcic2024towards . In these methods, natural language prompts are perturbed at various levels (e.g., character, word, sentence), and the functional correctness of code generated from original versus perturbed prompts is compared to measure performance degradation, regarding as the robustness measurement. Although these adversarial attacks can quantify the robustness, their reliability and generalization face challenges. Our observations indicate that different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed. Specifically, when we applied three adversarial attacks—character, word, and sentence levels respectively—used in recode wei2023towards to the MBPP Plus benchmark liu2024your , and instructed GPT-3.5-Turbo to generate results for both original and adversarial prompts. As a result, GPT-3.5-Turbo only performs less robustly under word-level adversarial attack with synonym replacement. These findings suggest that a single adversarial attack may not provide a holistic view of an LLM’s robustness in programming, and a suite of diverse attacks might be necessary for a comprehensive assessment of a single model. However, this approach complicates the comparative evaluation of robustness across different models. For instance, if one model demonstrates strong robustness against attack X but performs poorly against attack Y, while another model shows the opposite pattern, it becomes challenging to definitively determine which model is inherently more robust overall. This motivates a question: Can we overcome the configuration-dependency of adversarial attacks to achieve a unified evaluation of LLMs’ robustness in programming?
@@ 3 Evaluation
Benchmark Selection. EvaLoop is designed for versatility and seamless integration with existing programming benchmarks. In this study, we utilize two comprehensive benchmark suites, MBPP Plus liu2024your and MXEVAL athiwaratkun2022multi , which offer significant advancements over their predecessors. MBPP Plus enhances the original MBPP austin2021program by applying rigorous quality filtering to programming tasks and expanding the test case corpus 35-fold. MXEVAL extends MBPP’s Python-centric focus to encompass 12 distinct programming languages. For Framework 1 (Generation-Summarization Duality Loop), we evaluate the LLMs’ robustness in programming on 378 distinct tasks from MBPP Plus. The operational loop for this framework is limited to a maximum of 10 loops. For Framework 2 (Cyclical Code Translation Loop), we establish a circular code translation chain: Python → PHP → Ruby → JavaScript → Perl → Python. The four intermediate languages were randomly selected from the non-Python languages available in MXEVAL [^3] Our experimental results showed no significant differences between two such distinct random selections.. This cyclical translation process utilizes a carefully curated subset of 375 tasks from the multilingual MXEVAL.
@@
Examining ASL metric reveals another insight into LLMs’ robustness in programming. GPT-4.1 leads with an ASL of 3.42, closely followed by GPT-4.1-mini (3.40) and o3-mini (3.40), signifying not only strong initial accuracy but also remarkable performance consistency across sequential translation tasks, underscoring their robust nature and the superior. In stark contrast, open-source LLMs generally represent a relatively low ASL value, and LLaMA-3.1-8B records the lowest ASL with 2.11, whose trajectory shows a low baseline followed by continuous, pronounced degradation, highlighting difficulties in maintaining functional integrity. ASL is invaluable for a holistic assessment, moving beyond single-point evaluations to capture sustained operational efficacy and distinguish models with enduring performance integrity from those that falter under the cumulative stress of multi-step, cross-lingual code translations.
@@ 4 Related Work
LLMs’ Robustness in Programming. Code generation has emerged as the predominant practical application of AI in programming environments liang2024large , catalyzing extensive research into the robustness of LLMs in this domain. Empirical assessments of commercial systems reveal concerning vulnerabilities: GitHub Copilot produces inconsistent implementations for semantically equivalent prompts in 46% of cases, with correctness variations reaching 28% between prompt variants mastropaolo2023robustness . Similarly, Codex exhibits substantial sensitivity to minor linguistic perturbations in task specifications shirafuji2023exploring ; zhuo2023robustness . The research community has responded with LLMs’ robustness through diverse adversarial techniques chen2023evaluating ; zhang2024attacks ; jha2023codeattack ; bavcic2024jy61 alongside corresponding defensive strategies improta2025enhancing ; yang2024robustness ; ge2024demonstration . However, different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed, struggling in a unified evaluation. EvaLoop addresses this critical gap by leveraging the duality between generation and understanding, enabling unified robustness assessment across LLMs with the average number of sustainable loops, i.e., ASL . This novel approach provides a natural measure of LLMs’ robustness in programming.
Self-consistency in LLMs. Self-consistency in LLMs, originally introduced by Wang et al., represents a novel decoding strategy that significantly enhances reasoning accuracy by sampling diverse reasoning paths and selecting the most consistent answer through majority voting. Beyond using self-consistency to improve reasoning capability, researchers have begun exploring it as an evaluation metric for LLM performance min2023beyond ; lee2024evaluating .
IdentityChain min2023beyond is the most closely related work, employing a feedback loop between natural language specifications and code generation to assess semantic understanding in Code LLMs. However, IdentityChain fundamentally differs from EvaLoop in both objectives and methodology. While IdentityChain evaluates whether models can maintain semantic coherence in a single round-trip translation between natural language and code, EvaLoop addresses a distinct challenge: evaluating robustness through sustained iterative transformations that mirror real-world software evolution.
EvaLoop introduces several key innovations that distinguish it from IdentityChain: (1) a cross-language translation chain that stress-tests models across multiple programming paradigms rather than a single NL-to-code-to-NL cycle, (2) the Average Sustainable Loops (ASL) metric that quantifies long-term stability under continuous transformation pressure rather than one-time accuracy, and (3) a focus on functional correctness preservation through multiple iterations rather than semantic equivalence in a single translation. This fundamental difference in scope—from IdentityChain’s semantic consistency to EvaLoop’s robustness under iterative stress—makes EvaLoop uniquely suited for evaluating how Code LLMs perform in dynamic, evolving software development scenarios where code undergoes repeated modifications and translations.